# On the isomorphism problem for ultraproducts of $\mathrm{C}^{*}$-algebras in continuous model theory

### v1

@@ 5 Main Result
A noncommutative example is obtained by tensoring a unital noncommutative $\mathrm{C}^{\ast}$}-algebra with no non-trivial projections such as the Jiang-Su algebra $\mathcal{Z}$. Namely, for $X$ as above, $\mathcal{Z}\otimes C(X)$ is in $\mathcal{C}$.
Theorem 5.3.
[…]
Proof.
Thanks to section 3.3, we may assume $M$ is $\kappa$-saturated for large enough $\kappa$.
For $p,q\in\mathcal{P}(M)$, we define the relation $q\dashv p$ by
$q\dashv p\mathrel{\overset{\text{def}}{\iff}}pq\neq 0\land(1-p)q\neq 0\iff 1-\min\{\left\lVert{pq}\right\rVert,\left\lVert{(1-p)q}\right\rVert\}=0.$ (5.3.3)
We define $p_{\alpha},q_{\alpha}\in M$ and $M_{\alpha}\preceq M$ inductively for ordinals $\alpha\ (\leq\omega_{2})$. One can find a separable elementaly submodel $M_{0}$ of $M$ by section 2.4. Suppose that $M_{\alpha}$ is already defined. We take $p_{\alpha},q_{\alpha}\in\mathcal{P}(M)\setminus\{0,1\}$ such that the following conditions hold:
$\left\{\begin{array}[]{ll}q\dashv p_{\alpha}&\text{for all}\ q\in\mathcal{P}(M_{\alpha})\setminus\{0,1\},\\ q_{\alpha}\not\dashv p&\text{for all}\ p\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}.\end{array}\right$. (5.3.4)
To show existence of $p_{\alpha}\in\mathcal{P}(M)$, it is enough to check that the set
$t(x)\mathrel{:=}\{\left\lVert{x-x^{\ast}}\right\rVert,\left\lVert{x^{2}-x}\right\rVert\}\cup\left\{1-\min\{\left\lVert{xq}\right\rVert,\left\lVert{(1-x)q}\right\rVert\}:q\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\right\}$ (5.3.5)
is a type over $\mathcal{P}(M_{\alpha})\subseteq M$ because $M$ is $\kappa$-saturated. Let $n<\omega$ and $q_{i}\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\ (i<n)$. For each $\sigma\in 2^{n}$, we define
$q_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))q_{i}+\sigma(i)(1-q_{i})\right].$ (5.3.6)
If $q_{\sigma}\neq 0$ choose $p_{\sigma}\in\mathcal{P}(M)$ with $0\lneq p_{\sigma}\lneq q_{\sigma}$. Then, the sum $p\mathrel{:=}\sum_{\sigma}p_{\sigma}$ satisfies $p\neq 0,1$ and $q_{i}\dashv p$ for all $i<n$ since $q_{i}=\sum_{\sigma:\sigma(i)=0}q_{\sigma}$ and $1-q_{i}=\sum_{\sigma:\sigma(i)=1}q_{\sigma}$. Thus, $t(x)$ is finitely satisfiable in $M$. It follows from section 3.3 that $t(x)$ is a type over $\mathcal{P}(M_{\alpha})$.
Similarly, to take $q_{\alpha}\in\mathcal{P}(M)$, it is enough to show that the set
$t^{\prime}(x)\mathrel{:=}\{\left\lVert{x-x^{\ast}}\right\rVert,\left\lVert{x^{2}-x}\right\rVert\}\cup\left\{\min\{\left\lVert{px}\right\rVert,\left\lVert{(1-p)x}\right\rVert\}:p\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\right\}$ (5.3.7)
Let $n<\omega$ and $p_{i}\in\mathcal{P}(M_{\alpha})\setminus\{0,1\}\ (i<n)$. For each $\sigma\in 2^{n}$, we define
$p^{\prime}_{\sigma}\mathrel{:=}\prod_{i<n}\left[(1-\sigma(i))p_{i}+\sigma(i)(1-p_{i})\right].$ (5.3.8)
Since $1=\sum_{\sigma}p^{\prime}_{\sigma}$ there exist $\sigma^{\prime}\in 2^{n}$ such that $q\mathrel{:=}p^{\prime}_{\sigma^{\prime}}\neq 0,1$. Then, we have $q\not\dashv p_{i}$ for all $i<n$. Thus, $t^{\prime}(x)$ is finitely satisfiable in $M$. It follows from section 3.3 that $t^{\prime}(x)$ is a type over $\mathcal{P}(M_{\alpha})$.
By section 2.4, we take $M_{\alpha}\preceq M$ such that $M_{\alpha}\cup\{p_{\alpha},q_{\alpha}\}\subseteq M_{\alpha+1}$ and $\chi(M_{\alpha+1})\leq\chi(M_{\alpha})$. When $\gamma$ is limit ordinals, we define $M_{\gamma}\mathrel{:=}\overline{\bigcup_{\alpha<\gamma}M_{\alpha}}$.
Let $A\mathrel{:=}M_{\omega_{1}}$ and $B\mathrel{:=}M_{\omega_{2}}$. Then, $A,B\preceq M$ by section 2.4. Furthermore, $A$ and $B$ satisfy the following conditions respectively.
Claim 1.
Suppose that $\mathcal{U}\in\beta\omega\setminus\omega$. Then the following condition holds:
$\exists P\subseteq\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\}\mathrel{\text{s.t.}}\left\lvert{P}\right\rvert\leq\omega_{1}\land\left[\forall q\in\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\},\exists p\in P,q\dashv p\right].$ (5.3.9)
(Proof of Claim 1).
Let $P\mathrel{:=}\left\{[(p_{\alpha})_{n}]:\alpha<\omega_{1}\right\}$, $\left\lvert{P}\right\rvert\leq\omega_{1}$ is obviously. We show that $P$ satisfies the above condition. We note that $\mathcal{P}(\cdot)$ is commutative with any ultraproducts, since $\mathcal{P}(\cdot)$ is definable. Thus, let $q\in\mathcal{P}(A^{\mathcal{U}})\setminus\{0,1\}=\mathcal{P}(A)^{\mathcal{U}}\setminus\{0,1\}$, then we can take $q=[(q_{n})_{n}]$ where $q_{n}\in\mathcal{P}(A)\setminus\{0,1\}$ for $n<\omega$. By section 2.4, we have $A=\bigcup_{\alpha<\omega_{1}}M_{\alpha}$. Hence, for each $n<\omega$, we can take $\alpha_{n}<\omega_{1}$ such that $q_{n}\in\mathcal{P}(M_{\alpha_{n}})$. Let $\alpha\mathrel{:=}\sup_{n<\omega}\alpha_{n}$, then $\alpha<\omega_{1}$ because $\omega_{1}$ is regular. Thus, we put $p\mathrel{:=}[(p_{\alpha})_{n}]\in P$, then $q_{n}\dashv p_{\alpha_{n}}$ for all $n<\omega$. Therefore, by the section 3.1, we obtain $q\dashv p$. ∎
Claim 2.
Suppose that $\mathcal{V}\in\beta\omega\setminus\omega$. Then the following condition holds:
$\forall P\subseteq\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\},\left\lvert{P}\right\rvert\leq\omega_{1}\implies\left[\exists q\in\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\},\forall p\in P,q\not\dashv p\right].$ (5.3.10)
(Proof of Claim 2).
Let $P\subseteq\mathcal{P}(B^{\mathcal{V}})\setminus\{0,1\}$ with $\left\lvert{P}\right\rvert\leq\omega_{1}$ is given. Let $P=\left\{r^{\alpha}:\alpha<\omega_{1}\right\}$, and $r^{\alpha}=[(r^{\alpha}_{n})_{n}]$ where $r^{\alpha}_{n}\in\mathcal{P}(B)\setminus\{0,1\}\ (n<\omega)$ for each $\alpha<\omega_{1}$. By section 2.4, we have $B=\bigcup_{\beta<\omega_{2}}M_{\beta}$. Hence, for each $n<\omega$ and $\alpha<\omega_{1}$, we can take $\beta^{\alpha}_{n}<\omega_{2}$ such that $r^{\alpha}_{n}\in M_{\beta^{\alpha}_{n}}$. Let $\beta\mathrel{:=}\sup\left\{\beta^{\alpha}_{n}:n<\omega,\alpha<\omega_{1}\right\}$ ($\omega_{2}$, since $\omega_{2}$ is regular) and $q\mathrel{:=}[(q_{\beta})_{n}]$. Then, for any $\alpha<\omega_{1}$, $q_{\beta}\not\dashv r^{\alpha}_{n}$ for all $n<\omega$ and thus we obtain $q\not\dashv r^{\alpha}$, by section 3.1. ∎
From the above claims, it follows that $A^{\mathcal{U}}$ and $B^{\mathcal{V}}$ are not isomorphic for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$.
[…]
@@ 6 Conclusion
section 5 assumes commutativity for the set of all projections in a $\mathrm{C}^{\ast}$}-algebra, which is a quite strong condition. It is natural to ask whether the same result holds for more general $\mathrm{C}^{\ast}$}-algebras.
Question 6.1.
[…]
We remark that, in general, when $\mathrm{T}$ is a stable $\mathcal{L}$-theory, it is known that any ultrapowers of $\mathcal{L}$-structures with density character $\leq{\mathfrak{c}}$ are isomorphic in ZFC ([FHS2:2014, Theorem 5.6 (1)]). Therefore, this question is negated when $\mathrm{T}$ is stable. However, any theory of infinite-dimensional unital $\mathrm{C}^{\ast}$}-algebras is unstable (see [FHS1:2013, Lemma 5.3] and [FHS2:2014, Theorem 5.5]). So, [FHS2:2014, Theorem 5.6 (1)] cannot be applied to this question.
Question 6.2.
[…]
To approach this problem, we would like to investigate the situation of section 5.
@@
Therefore, by applying section 5 to $M=C(2^{\omega})$, we obtain the following result:
@@
[…]
@@
Suppose that CH fails. By section 5 with $M=C(2^{\omega})$, there exist $A,B\equiv M$ such that $\chi(A),\chi(B)\leq{\mathfrak{c}}$ and $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$. By section 3.1, there exist a set $I$ and non-principal ultrafilters $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $A^{\mathcal{U}}\cong B^{\mathcal{V}}$.
@@
Thus, section 6 for abelian $\mathrm{C}^{\ast}$}-algebras can be reformulated as the following question about topological spaces:
@@
[…]

### v2

@@ 5 Main Result
A noncommutative example is obtained by tensoring a unital noncommutative $\mathrm{C}^{\ast}$}-algebra with no non-trivial projections such as the Jiang-Su algebra $\mathcal{Z}$. Namely, for $X$ as above, the theory of $\mathcal{Z}\otimes C(X)$ satisfies the conditions in section 5.
@@ 6 Conclusion
In conclusion, we discuss a topological applications of section 5 and pose some questions. We begin by considering the commutative situation of section 5.
@@
Therefore, by applying section 5 to $C(2^{\omega})$, we obtain the following result:
@@
Under the failure of the Continuum Hypothesis, there exist Stone spaces $X$ and $Y$ satisfying the following properties:
@@
Suppose that the Continuum Hypothesis fails. By section 5 with $M=C(2^{\omega})$, there exist $A,B\equiv M$ such that $\chi(A),\chi(B)\leq{\mathfrak{c}}$ and $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$. By section 3.1, there exist a set $I$ and non-principal ultrafilters $\mathcal{U},\mathcal{V}\in\beta I\setminus I$ such that $A^{\mathcal{U}}\cong B^{\mathcal{V}}$.
@@
We now turn to a discussion of generalizations of section 5 and section 5. We note that if a theory $\mathrm{T}$ is stable (in the sense of continuous model theory; cf. [FHS2:2014, Definition 5.2]), then any ultrapowers of model of $\mathrm{T}$ with density character at most $\mathfrak{c}$} are isomorphic, without assuming the Continuum Hypothesis ([FHS2:2014, Theorem 5.6 (1)]). Thus, we consider the following question:
Question 6.5.
Assume that the Continuum Hypothesis fails. Let $\mathrm{T}$ be a complete unstable theory with $\chi(\mathcal{L},\mathrm{T})\leq\aleph_{0}$. Do there exist $A,B\models\mathrm{T}$ with $\chi(A),\chi(B)\leq{\mathfrak{c}}$ such that $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$?
We note that the assumptions of section 5 immediately imply that the theory $\mathrm{T}$ has the order property (cf. [FHS2:2014, Definition 5.2]), which is equivalent to $\mathrm{T}$ being unstable (see [FHS2:2014, Theorem 5.5]). It is known that any theory of a unital infinite-dimensional $\mathrm{C}^{\ast}$}-algebras have the order property (see [FHS1:2013, Lemma 5.3]). Therefore, as a special case of the above question, the following question is also of interest.
Question 6.6.
Assume that the Continuum Hypothesis fails. Let $\mathrm{T}$ be the theory of a unital infinite-dimensional $\mathrm{C}^{\ast}$}-algebra. Do there exist $A,B\models\mathrm{T}$ with $\chi(A),\chi(B)\leq{\mathfrak{c}}$ such that $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$?
@@
Assume that the Continuum Hypothesis fails. Does there exist a complete theory $\mathrm{T}$ of $\mathrm{C}^{\ast}$}-algebras with separable models $A,B\models\mathrm{T}$ such that $A^{\mathcal{U}}\not\cong B^{\mathcal{V}}$ for all $\mathcal{U},\mathcal{V}\in\beta\omega\setminus\omega$?
In particular, by the discussion of section 6, the abelian $\mathrm{C}^{\ast}$}-algebra case can be reformulated the following question:
Question 6.8.
Assume that the Continuum Hypothesis fails. Do there exist Stone spaces $X$ and $Y$ satisfying the following properties?